There are approximately 331 million people in the United States. It is estimated that 1 out of every 700 people in the US have polycystic kidney disease. That means that 699 out of every 700 people are healthy. Answer the following questions using this information, your answers to the previous question, and the Hardy-Weinberg equation. Show your work and include the variable/s you are solving for in each question.

a. What is the frequency of healthy people? (Round to the nearest four decimal points)

b. What is the frequency of the healthy allele? (Round to the nearest four decimal points)

c. What is the frequency of the allele that causes polycystic kidney disease? (Round to the nearest four decimal points)

d. How many (rounded to the nearest whole person) people in the US have each of the genotypes that causes polycystic kidney disease?
Answer #1

a. frequency of healthy people :

699 are healthy out of 700.

Frequency of healthy people = 699 / 700 = 0.9986

b. Frequency of people suffering from polycystic kidney disease:

1 out 700 are suffering.

Frequency = 1 / 700 = 0.0014

c. Frequency of the allele causing polycystic kidney disease:

Square root of people suffering from kidney disease. because each individuals suffering will have 2 alleles for the disease.

Frequency of the allele = 0.0378

d.People suffering from kidney disease are --

frequency of the genotype x total population

= 0.0014 x 331 x 10^6 = 463400 people or 0.4634 millions

People normal in the population:

Frequency of people who are normal x population

0.9986 x 331 x 10^6 = 330536600 persons. or 330.5366 millions.